# HREB-CRF: Hierarchical Reduced-bias EMA for Chinese Named Entity Recognition

### v1

@@ 4.6 Case Analysis
In Example-1 of Figure-5, the performance of the model trained on MSRA dataset is verified using normal press releases. The results show that proposed model can effectively extract entity information for regular text. For the baseline model, it recognizes the text of the school-college combination as two different entities. When considering ground truth, it is generally believed that school+college is a complete entity.
@@
In Example-2 of Figure-7, an extreme scenario was constructed. The same Chinese character was expressed as a noun, a verb, and an adjective simultaneously. When these were combined into a sentence, it was verified whether the model could output correctly in extreme cases. The experimental results indicate that the two protagonists could be effectively identified by the proposed model. However, for the baseline model, although the two protagonists were covered, the entity boundary was found to be unclear.
@@
The analysis presented in Figure-6 addresses two primary challenges in entity recognition, as outlined in the introduction: 1) ambiguity in boundary identification (example-3); 2) errors in entity recognition (example-4). A comparative evaluation demonstrates that the proposed model effectively resolves these issues in the selected examples, exhibiting superior performance relative to the baseline model.
4.7 Analysis and Discussion
Through comparative experiments and ablation studies, an in-depth analysis of various aspects of model performance was conducted. The following are the main findings and discussions derived from the experimental results:
1) Proposed method can be matched with the comparative experiment and the ablation experiment to prove that the model can adapt to different types of semantic information under different scene tasks and adaptively improve the robustness of CNER. Comparing the experimental results of Table-2, Table-4, and Table-3, proposed model achieves the best performance in both regression and F1 indicators. The ablation experiment in Table-5 shows that the hierarchical exponential attention and reduced-biased modules can effectively expand the deep dimensions of the data compared with the baseline model. It performs well in the conventional tasks in Figure-5.
2) Proposed research shows that the Hierarchical Reduced-bias EMA architecture can bring accurate feature extraction for CNER tasks. The modeling of the attention mechanism is to understand the text content from the vector dimension, which is used in the latest LLMs models (such as GPT, Llama). In the field of transfer learning, inspired by DAE[27], proposed model can be used in other sequence modeling tasks. The sequence modeling task uses BERT to model the text vectorization input, and the output is the probability weight of the same length as the input text. It is believed that the model can be applied to NER tasks such as text understanding, text classification, and emotion recognition simultaneously.
3) Increasing the number of attention heads will significantly increase the number of parameters in the multi-head attention mechanism, thereby increasing the training time and inference time of the model. When selecting the number of attention heads, proposed study comprehensively considered the trade-off between performance and efficiency and selected EMA head of 64.
5 Conclusion
@@ Acknowledgments
This was was supported in part by Agency for Science, Technology and Research.
Appendix
5.1 CRF Decoding
CRF was introduced to strengthen the shortcomings of HMM in feature selection and dependency assumptions. HMM is a generative model that models the joint probability $P(x,y)$, that is, the joint distribution of the observed data and the label sequence, expressed as the Equation-22. This requires HMM to make strong assumptions about the relationship between the observed data and the hidden state (such as the independent and identically distributed assumption and the Markov assumption). In contrast, CRF is a discriminative model that directly models the conditional probability $P(y|x)$ without making independence assumptions on the observed data. Therefore, CRF introduces feature functions and weights into the model and maximizes the conditional probability of given observed data by learning these parameters. It is able to directly model the label sequence without assuming the Markov property between the observed data and the hidden state. This enables CRF to better capture long-distance dependencies in the sequence and automatically learn the relationship between features without manually designing features.
$P(x,y)=P(y_{1})\prod_{t=2}^{T}P(y_{t}|y_{t-1})\prod_{t=1}^{T}P(x_{t}|y_{t})$ (22)
The calculation process of CRF involves constructing and optimizing the conditional probability $P(y|x)$, that is, the probability of predicting the label sequence $y$ given the observation sequence $x$. The core of CRF is to maximize this conditional probability by defining feature functions and learning weights. Specifically, the derivation formula of CRF is:
$P(y|x)=\frac{1}{Z(x)}\exp\left(\sum_{t=1}^{T}\sum_{k}\lambda_{k}f_{k}(y_{t-1},y_{t},x,t)\right)$ (23)
where $f_{k}(y_{t-1},y_{t},x,t)$ is the feature function that defines the relationship between the label sequence $y$ and the observation sequence $x$. $\lambda_{k}$ is the weight parameter of the feature function, which is learned through training. $Z(X)$ is the normalization term (partition function) that ensures that the sum of the conditional probabilities is the Equation-24.
$Z(x)=\sum_{y}\exp\left(\sum_{t=1}^{T}\sum_{k}\lambda_{k}f_{k}(y_{t-1},y_{t},x,t)\right)$ (24)
5.2 Model Training Performance
Figure-8, Figure-9 and Figure-10 are the training processes of MSRA, Weibo and Resume respectively. The training process shows that MSRA reaches asymptotic performance at about 60 steps, Weibo reaches asymptotic performance at 15 steps, and Resume reaches asymptotic performance at 85 steps.
Figure 8: Training progress metrics for MSRA.
Figure 9: Training progress metrics for Weibo.
Figure 10: Training progress metrics for Resume.

### v2

@@ IV-F Case Analysis
In Example-1 of Figure 4, the performance of the model trained on MSRA dataset is verified using normal press releases. The results show that proposed model can effectively extract entity information for regular text. For the baseline model, it recognizes the text of the school-college combination as two different entities. When considering ground truth, it is generally believed that school+college is a complete entity.
@@
In Example-2 of Figure 6, an extreme scenario was constructed. The same Chinese character was expressed as a noun, a verb, and an adjective simultaneously. When these were combined into a sentence, it was verified whether the model could output correctly in extreme cases. The experimental results indicate that the two protagonists could be effectively identified by the proposed model. However, for the baseline model, although the two protagonists were covered, the entity boundary was found to be unclear.
@@
The analysis presented in Figure 5 addresses two primary challenges in entity recognition, as outlined in the introduction: 1) ambiguity in boundary identification (example-3); 2) errors in entity recognition (example-4). A comparative evaluation demonstrates that the proposed model effectively resolves these issues in the selected examples, exhibiting superior performance relative to the baseline model.
IV-G Analysis and Discussion
Through comparative and ablation studies, key findings on model performance are summarized:
1) Proposed model adapts to diverse semantic information across scene tasks, enhancing CNER robustness. As shown in Tables II, IV and III, proposed method achieves top performance in regression and F1 scores. Ablation studies in Table V confirm that hierarchical exponential attention and reduced-bias modules expand data depth, excelling in conventional tasks in Figure 4 and domain tasks in Figures 6 and 5.
2) The Hierarchical Reduced-bias EMA architecture enables precise CNER feature extraction. Inspired by DAE [27], the model extends to sequence tasks (e.g., text classification, emotion recognition) using BERT-based vectorization.
3) Increasing attention heads raises parameters, prolonging training/inference time. In order to balanceto balance performance and efficiency, 64 EMA heads were selected.
V Conclusion
@@ Acknowledgments
This research was funded by Programme MARS (Programme of Maritime AI Research in Singapore) with funding grant number SMI-2022-MTP-06 by Singapore Maritime Institute (SMI).